there are four numbers such that each is the sum of the previous number and 4. Find the numbers if their sum is 64.
Answer by addingup(3677)

We'll call the first number x:
x+x+4+x+4+4+x+4+4+4= 64
Add all the x and all the 4 (or, for the 4, you can multiply 4*4):
4x+24= 64 Subtract 24, both sides
4x= 40 Divide both sides by 4
x= 10
Proof:
10+(10+4)+(10+4+4)+(10+4+4+4)= 64
10+14+18+22= 64 These are your numbers. Add them up and see proof:
64= 64 we have the correct answer